OpenStudy (anonymous):

Consider the function f(x) = x + 2/x, defined over the domain 1 ≤ x ≤ 3. Draw its graph by following the steps: a) Find f′(x); b) Solve f′(x) = 0 c) Find where the function is increasing and decreasing d) Find f′′(x) e) Find where the function is concave up and concave down f) Find the coordinates of all points of interest

OpenStudy (anonymous):

I already have part a and b but im stuck on c

OpenStudy (mertsj):

The function is a line. It has positive slope. It is increasing over all intervals.

OpenStudy (anonymous):

is there a way i can prove that mathematically?

OpenStudy (mertsj):

Draw the graph

OpenStudy (anonymous):

awesome thanks. i can do part d but what about part e

OpenStudy (mertsj):

Also if the first derivative is positive, the function is increasing.

OpenStudy (mertsj):

If the second derivative is positive, it is concave up

OpenStudy (anonymous):

the whole function or only some of it?

OpenStudy (mertsj):

On the interval where it is positive

OpenStudy (mertsj):

f(x) is concave up on I iff f''(x) \>= 0$ on I.

OpenStudy (anonymous):

and to find the coordinates in part f i would use the original function and only include the numbers that are within the domino right?

OpenStudy (anonymous):

*domain

OpenStudy (anonymous):

do i just make x+(2/x)=0 and find values for x?

OpenStudy (mertsj):

Where is x + (2/x)=0 coming from?

OpenStudy (anonymous):

the orig function

OpenStudy (anonymous):

is that the right way to do it. i dont know how to find the intersect if its a straight line??

OpenStudy (mertsj):

That's not the original function you posted.

OpenStudy (anonymous):

dang typo. it was meant to be F(x)=x+(2/x)

OpenStudy (anonymous):

i did part a and be right and the fist derivative is positive so it is increasing. the second derivative is positive also so it is concave up so what steps do i need to take for finding the coordinates of all points of intersect

OpenStudy (zarkon):

What did you get for \(F'(x)\)?

OpenStudy (anonymous):

1-(2/x^2)

OpenStudy (zarkon):

ok...so on what interval is \(F\) increasing?

OpenStudy (anonymous):

Mertsj said it was increasing over all intervals

OpenStudy (zarkon):

that was your previous function

OpenStudy (anonymous):

ohhhh so is it sill increasing?

OpenStudy (zarkon):

you tell me

OpenStudy (anonymous):

ok so if the derivative is f'(x)=1-(2/x2), then would it mean that it decreases because the x in the equation is negative? ?

OpenStudy (anonymous):

Ive worked out hat for part b) when f'(x) = 0 , x=plus and minus the square root of 2. Is that right?